There is a 0.9968 probability that a randomly selected 50-year-old female lives through the year (based on data from the U.S. Department of Health and Human Services). A Fidelity life insurance company charges $226 for insuring that the female will live through the year. If she does not survive the year, the policy pays out $50,000 as a death benefit.
From the perspective of the 50-year-old female, what are the values corresponding to the two events of surviving the year and not surviving?
If a 50-year-old female purchases the policy, what is her expected value?
WORK TRIED:
In the event she lives, the value is -$226. In the event she dies, the value is $49,774.

Answer by stanbon(75887)

There is a 0.9968 probability that a randomly selected 50-year-old female lives through the year (based on data from the U.S. Department of Health and Human Services).
-------------------
A Fidelity life insurance company charges $226 for insuring that the female will live through the year. If she does not survive the year, the policy pays out $50,000 as a death benefit.
From the perspective of the 50-year-old female, what are the values corresponding to the two events of surviving the year and not surviving?
----
Ans: -226 ; 50,000-226 = 49774
-------------------------
If a 50-year-old female purchases the policy, what is her expected value?
WORK TRIED:
In the event she lives, the value is -$226. In the event she dies, the value is $49,774.
----
E(x) = 0.9968*(-226) + 0.0032(49774) = -$66
==================================================
Cheers,
Stan H.
===========